OpenStudy (anonymous):

Solve the expression T(m-n)/s=r A. T=m-n/rs B. T=rs(m-n) C. T=rs/m-n D. T=r+s/m-n

OpenStudy (anonymous):

T(m-n)/s=r ............multiply both sides by r T(m-n)=rs .................... divide both sides by (m-n) T=rs/(m-n)

OpenStudy (anonymous):

technically C but for it to agree with bedmas the m-n should include brackets i.e. (m-n)

OpenStudy (anonymous):

Oh thats right both sides by m-n ok I got it thanks bro

OpenStudy (anonymous):

and by r i meant s in the first step

OpenStudy (anonymous):

ok I gotcha thanks